a person invests $20000 partly at 5%, partly at 6% and partly at 6.5%. the total interest is $1170. three times of the amount invested at 6% equals the amount invested at 5% and 6.5% combined. how much is invested in each?
Answer by stanbon(75887)

a person invests $20000 partly at 5%, partly at 6% and partly at 6.5%. the total interest is $1170. three times of the amount invested at 6% equals the amount invested at 5% and 6.5% combined. how much is invested in each?
-------------------------------
Equations:
x + y + z = 20000
0.05x + 0.06y + 0.065z = 1170
3y = x + z
---------------------------
Modify and Rearrange:
x + y + z = 20000
50x+60y+65z=1170000
x - 3y + z = 0
---------------------------
Using a matrix method I got the following.
x = 7000
y = 5000
z = 8000
----
That answer does make sense. In my earlier post
I had dropped a zero on the 1170*1000
===============================================================
Cheers,
Stan H.
-------------------